OpenStudy (anonymous):

Does anyone take Geometry A at CCA? I need help with Unit 3 Lesson 10.

OpenStudy (idku):

You can ask a question directly, because this type of questions is reportable.

OpenStudy (anonymous):

Which of the following equations equations would graph a line parallel to 3x-y=7? A. y=7x-3 B. y=-3x+7 C. y=-7x+3 D. y=3x

OpenStudy (idku):

the parallel line must have the same slope.

OpenStudy (idku):

First write 3x-y=7 in a y=mx+b from.

OpenStudy (anonymous):

y=3x+7?

OpenStudy (idku):

yes, but negative 7. (doesn't matter in this case, really, but still...) so which of the choices has the same slope?

OpenStudy (anonymous):

C?

OpenStudy (idku):

No. y=mx+b, m is the slope.

OpenStudy (anonymous):

B?

OpenStudy (idku):

Why are you guessing?

OpenStudy (anonymous):

Because I'm confused

OpenStudy (idku):

y=3x-7 what other equation also has the same slope. (slope =3)

OpenStudy (anonymous):

D

OpenStudy (idku):

Yes, but do you understand how?

OpenStudy (anonymous):

yes

OpenStudy (idku):

Good!